Below are nominal GDP and GDP deflators for four years.

Year

Nominal GDP in billions GDP Deflator

Year 1

$19,000 100

Year 2

19,560 102

Year 3

20,495 105

Year 4

21,280 106


a. Calculate real GDP in each year.

Instructions: Round your responses to the nearest whole number.

    For Year 1, GDP is $ billion.

    For Year 2, GDP is $ billion.

    For Year 3, GDP is $ billion.

    For Year 4, GDP is $ billion.

b. Did the percentage change in nominal GDP exceed the percentage change in real GDP in any of the last three years listed?

    Year 2: (Click to select)NoYes.

    Year 3: (Click to select)YesNo.

    Year 4: (Click to select)NoYes.

c. In which year did society’s welfare increase the most?

    (Click to select)It's impossible to say[A(3)][A(1)][A(2)][A(4)].
Answer #1

a)

Nominal GDP Real GDP = GDP deflator GDP deflatox 100

Real GDP year 1 = 19,000

19,560 Real GDP year 2 = 102 x 100

Real GDP year 2 = 19,177

20,495 Real GDP year 3 = 105 X 100

Real GDP year 3 = 19,519

21,280 Real GDP year 4 = 106 x 100

Real GDP year 4 = 20,076

--------------------------------------------------------------------------------------------

b)

19560 - 19000 % change in nominal GDP year 2 = - 19000 X 100

% change in nominal GDP year 2 = 2.95%

% change in real GDP year 2 = 19177 - 19000 19000 -X 100

% change in real GDP year 2 = 0.93%

20495 - 19560 % change in nominal GDP year 3 = X 100 19560

% change in nominal GDP year 3 = 4.78%

19519 - 19177 % change in real GDP year 3 = = 19177 - X 100

% change in real GDP year 3 = 1.78%

% change in nominal GDP year 4 = 21280 - 20495 20495 --x 100

% change in nominal GDP year 4 = 3.83%

% change in real GDP year 4 = 20076 - 19519 19519 -X 100

% change in Real GDP year 4 = 2.85%

-------------------------------------------------------------------------------------

c)

In year 4, the society's welfare increased the most because the % change in real GDP in year 4 is the highest compared to year 3 and year 2.